Circular No. 5662

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1992 UY_{4}

C. S. Shoemaker reports her measurements of an asteroidal object found on films taken in late October by E. M. Shoemaker, H. E. Holt, D. H. Levy and herself with the 0.46-m Palomar Schmidt. Computations by the undersigned indicated an Amor-like orbit with a perihelion distance very close to 1 AU. In collaboration with S. Nakano (Sumoto, Japan), further observations were obtained in November by A. Sugie (Dynic Astronomical Observatory) and T. Seki (Geisei).

1992 UT α_{2000} δ_{2000} Mag. Observer
Oct. 25.48541 4^{h}04^{m}09.^{s}60 +36^{o}16′18′′.6  17.2 Shoemaker
25.52152 4 04 06.49 +36 15 35.2  _{”}
26.44166 4 03 02.60 +35 58 26.7  _{”}
26.49305 4 02 58.19 +35 57 24.9  _{”}
Nov. 17.58125 3 38 52.61 +29 49 13.4  17.0 Sugie
17.59514 3 38 51.96 +29 49 02.5  _{”}
18.56319 3 38 04.68 +29 35 14.3  _{”}
18.57569 3 38 04.15 +29 35 06.4  _{”}
23.46910 3 34 36.58 +28 29 15.9 17 Seki
23.48785 3 34 35.92 +28 29 07.0  _{”}
25.50139 3 33 26.23 +28 04 01.2 17 _{”}
25.51285 3 33 25.77 +28 03 54.4  _{”}

Nakano supplies elements based on 11 of the above positions:

T = 1992 Sept. 8.852 TT  ω = 37.^{o}814
e = 0.62836  Ω = 308.684 }2000.0
q = 1.00919 AU  i = 2.874

a = 2.71552 AU n^{o} = 0.220255 P = 4.47 years These elements indicate opposition in June 1992 (δ = −29^{o}) and an

approach to within 0.031 AU of the earth on Sept. 7.9 (δ = +81^{o}).

1992/93 α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Nov. 24 3^{h}34.^{m}27 +28^{o}22.^{′}7 0.422 1.405 170.^{o}3 6.^{o}8 16.6
Dec. 4 3 30.35 +26 31.7 0.517 1.487 163.0 11.2 17.3
14 3 30.32 +25 10.8 0.628 1.571 153.4 16.3 18.0
24 3 33.70 +24 16.7 0.755 1.655 144.1 20.4 18.7
Jan. 3 3 39.97 +23 44.9 0.897 1.740 135.4 23.4 19.3

1992 November 27 Gareth V. Williams